After a 13% discount, a new DVD player was selling for $217.50. What was the original price of the DVD player?
Answer by jim_thompson5910(35256)

x = original price

x - (13% of x) = 217.50

x - 0.13x = 217.50

0.87x = 217.50

x = 217.50/0.87

x = 250

The original price was $250